You have a money jar containing nickels(.05) and quaters(.25) worth $1.55. The money jar contains 11 coins. How many of each coin do you have?
Answer by matthew_sessoms(39)

x=# of nickles
y=# of quarters


We'll have a system of equations...
x + y = 11
.05x + .25y = 1.55


I'll use substitution. Solve for "x" in equation #1.
x = 11 - y
Now, plug this in into the second equation...


.05(11-y)+.25y=1.55
.55-.05y+.25y=1.55
.2y=1.00
y=5


Now, plug y=5 into either of the orginal equations & solve for "x". I chosed equation 1 because it looks the easiest.


x+5=11
x=6


6 nickels ($0.30)
5 quarters ($1.25)

Thanks
MS